# Flexible Process Variant Binding in Information Systems with Software Product Line Engineering

### v1

@@ 7.3 Software Product Line Engineering for Variable Business Processes
The concept of SPL Engineering has been applied to business processes in other works such as [50],[1], [43], [9], and [47].
@@
In [50], the authors introduce means to mark elements in a process model as variation points. Different variants as well as their binding times may be specified as realisation for these variation points. A proof of concept is built on top of the Eclipse plugin architecture. However, the work focuses on the modelling aspect of variation points in process models rather than the implementations details of variation points. The work lacks a detailed description of the variability mechanisms used to bind the variants and a description of the management of feature interactions. In our approach, we present an in-depth approach on how to use existing variability mechanisms to bind features at compile time, start time, and runtime. Furthermore, our approach describes how to handle feature interactions when applying concepts of SPL Engineering to variable business processes.
@@
SPL Engineering is used in [1] as well to compose business processes. The authors argue that the services called by a business process may be owned by a third-party vendor and may be highly configurable. In order to compose a business process comprising services those services are each treated as an SPL whereas the business process itself is considered a multiple SPL. The authors focus on aggregating the feature models of the SPLs, automatically checking the feature selection against the aggregated feature model, and determining which services are able to accept the output data of other services as input data. Although the authors describe how to check whether or not the services are compatible (e.g. in regard to data structures), they fail to describe how the services are automatically composed, what technical mechanism shall be used to automatically bind and connect the services. Furthermore, dynamic binding of the services and multiple binding for one activity are neglected by this work. Besides, the considered business processes only include automated activities and do not contain user tasks.
@@
Concepts of SPL like feature modelling have been applied to cross-organisation business processes in [43]. However, the approach is not holistic and does not describe how to systematically and automatically instantiate software products derived from the SPL.
@@
In [9], the customers’ requirements and the corresponding parts of the business process that implement these requirements are mapped in a decision table, which can be used to configure and derive a business process from the product line. However, the authors do not describe how the configuration works from a technical perspective.
@@
SPL Engineering concepts were also applied to automatically generate process model variants. Changes in process models can be automatically propagated to the derived process model variants [47]. Our work distinguishes itself from the above one in that it focuses on the automatic generation of PAISs and not only on the automatic generation of process models.
@@ 8 Conclusions
The presented work complemented the approach of PAIS Product Lines towards more flexibility. A plugin mechanism was introduced that is based on FeatureHouse and parametrization. Using this plugin mechanism, activity implementations may be bound at compile time, start time, and runtime. Furthermore, FeatureHouse was used to compose the process data structure. Different organisations may require varying process data structures for the same business process. By composing the process data structure using FeatureHouse, the process data structure may be adapted to an organisation’s specific needs during PAIS Product derivation. In addition, the plugin mechanism enables the selection if multiple implementations for an activity. It has been deduced that aggregation code is required when selecting multiple implementations for one activity that write to the same part of the process data structure to prevent the implementations from overwriting each others data. Aggregation code is additional code being included when multiple implementations for one activity are selected and that aggregates the output data of those implementations. Finally, FeatureIDE is used as tool support for the phases domain analysis, domain implementation, requirements analysis, and product derivation.
@@
The presented approach has limitations, which will be addressed in future research. It focuses on varying activity implementations and assumes a fixed and immutable control flow of the process model during PAIS Product derivation. Future research shall reveal how the control flow may be adapted during PAIS Product derivation. The approaches of base and reference process models might be combined and included into the approach of PAIS Product Lines. Besides, only activities may constitute a variation point, i.e only the implementation of activities may be selected and substituted. Future research shall outline how events may constitute a variation point and how their implementation may be substituted.

### v2

@@ 8.3 Software Product Line Engineering for Variable Business Processes
The concept of SPL Engineering has been applied to business processes in other works such as Taffo Tiam et al. (2015),Acher et al. (2012), de S. Gimenes et al. (2008), Boffoli et al. (2008), and Sinnhofer et al. (2015).
@@
In (Taffo Tiam et al., 2015), the authors introduce means to mark elements in a process model as variation points. Different variants as well as their binding times may be specified as realisation for these variation points. A proof-of-concept is built on top of the Eclipse plugin architecture. However, the work focuses on the modelling aspect of variation points in process models rather than the implementations details of variation points. The work lacks a detailed description of the variability mechanisms used to bind the variants and a description of the management of feature interactions. In our approach, we present an in-depth approach on how to use existing variability mechanisms to bind features at compile time, start time, and runtime. Furthermore, our approach describes how to handle feature interactions when applying concepts of SPL Engineering to variable business processes.
@@
SPL Engineering is used by Acher et al. (2012) as well to compose business processes. The authors argue that the services called by a business process may be owned by a third-party vendor and may be highly configurable. In order to compose a business process comprising services those services are each treated as an SPL whereas the business process itself is considered a multiple SPL. The authors focus on aggregating the feature models of the SPLs, automatically checking the feature selection against the aggregated feature model, and determining which services are able to accept the output data of other services as input data. Although the authors describe how to check whether or not the services are compatible (e.g. in regard to data structures), they fail to describe how the services are automatically composed, what technical mechanism shall be used to automatically bind and connect the services. Furthermore, dynamic binding of the services and multiple binding for one activity are neglected by this work. Besides, the considered business processes only include automated activities and do not contain user tasks.
@@
Concepts of SPL Engineering like feature modelling have been applied to cross-organisation business processes by de S. Gimenes et al. (2008). However, the approach is not holistic and does not describe how to systematically and automatically instantiate software products derived from the SPL.
@@
In (Boffoli et al., 2008), the customers’ requirements and the corresponding parts of the business process that implement these requirements are mapped in a decision table, which can be used to configure and derive a business process from the product line. However, the authors do not describe how the configuration works from a technical perspective.
@@
SPL Engineering concepts were also applied to automatically generate process model variants. Changes in process models can be automatically propagated to the derived process model variants (Sinnhofer et al., 2015). Our work distinguishes itself from the above one in that it focuses on the automatic generation of PAISs and not only on the automatic generation of process models.
8.4 Software Process Lines
SPL Engineering concepts have also been applied to software processes (Nogueira Teixeira et al., 2019), which define the steps to develop software products (i.e. the process that describes how software development teams collaborate). SPEM [^15] https://www.omg.org/spec/SPEM/2.0/PDF is a notation to define and illustrate software processes. As SPEM models cannot be executed, there have been approaches to convert SPEM models to BPMN models in order to execute them, e.g. Dashbalbar et al. (2017) manually translate a software process in the context of software testing modelled in SPEM to BPMN.
Agh and Ramsin (2021) have analysed Scrum [^16] https://scrumguides.org/docs/scrumguide/v2020/2020-Scrum-Guide-US.pdf, which is an agile approach commonly employed in software processes. The authors identified the variation points in Scrum, which include roles, e.g. the Product Owner role may be exercised by an analyst or a project manager (both require domain knowledge of the target product). Other variation points are the length of the development cycles (i.e. sprint length) and the technique for estimating the bearable workload for the next development cycles (i.e. sprint velocity). The Scrum process is modelled using the notation SPEM. Transformation rules are then automatically applied to the process model resulting in variants, whereby the resulting process models are manually adjusted. In a case study, Scrum teams derived process variants according to their needs and adopted it in their daily work. The authors argue their customisable Scrum software process contains best practices from which software development teams can select and adopt their best fit in order to address shortcomings in their current Scrum processes. As the approach addresses a software process modelled in SPEM, the process model cannot be executed.
Aleixo et al. (2010) introduced an approach that allows modelling a software process using the Unified Method Architecture (UMA), which is an evolution of SPEM 1.1. Variability is represented as a feature model. After deriving a process variant, the process variant can be transformed to a workflow specification that can be executed by the WfMS jBPM [^17] https://docs.jbpm.org/latest/jbpm-docs/html_single/. All the activities of workflow specification are implemented as forms (i.e. user tasks). While jBPM now uses BPMN as modelling notation to execute process models, at the time of the paper, the BPMN standard did not support native execution of process models (Geiger et al., 2018).
Garcia et al. (2015) pictured a software process using UMA for the development of service-oriented software products which can be tailored to the needs of a project. The software process is first modelled incorporating variability. Then concrete software processes may be derived by selecting the required elements at the variation points with respect to the needs of the project. Finally, using a custom tool the software process model is converted to a BPMN model which is manually adjusted. A WfMS can execute the converted BPMN model whose activities are all implemented by a standard form (i.e. user task). While the converted process can be executed, the variability occurs solely in the process model. There is no variability on the implementation level. In addition, it is a very simple process consisting of only user tasks, which all comprise the same standard form.
The literature review conducted by Nogueira Teixeira et al. (2019) revealed that most of the researchers applying SPL Engineering concepts to software processes use SPEM as a modelling notation for the control flow and only a few approaches (the two described above) allow executing processes. In contrast to customisable process models, software process lines allow customising elements of the process (e.g. activities, roles) and rarely the control flow (Nogueira Teixeira et al., 2019). Finally, the authors point out that software process lines might benefit from further incorporating business process approaches, which allow executing the software processes.
In summary, the approach of software process lines focuses on variability in software processes whereas our approach addresses business processes in general, which has the following implications: Software process lines include variation points that need to be implemented organizationally and which is not a technical challenge (e.g. staffing of a role). Although some approaches allow converting process models to BPMN for execution, mostly the SPEM standard is used to model software processes, which lack the ability to be executed. Consequently, software process lines only allow deriving process variants on an illustrative level. In conclusion, our approach does not include software processes but business processes in general and allows automatically deriving entire information systems implementing business processes (i.e. PAIS Products) by solving variation points on an implementation level whereas software process lines focus on variability on the model level.
@@ 9 Conclusions
The presented work complemented the approach of PAIS Product Lines towards more flexibility. A plugin mechanism was introduced that is based on FeatureHouse and parametrization. Using this plugin mechanism, activity implementations may be bound at compile time, start time, and runtime. Furthermore, FeatureHouse was used to compose the process data structure. Different organisations may require varying process data structures for the same business process. By composing the process data structure using FeatureHouse, the process data structure may be adapted to an organisation’s specific needs during PAIS Product derivation. In addition, the plugin mechanism enables the selection if multiple implementations for an activity. It has been deduced that aggregation code is required when selecting multiple implementations for one activity that write to the same part of the process data structure to prevent the implementations from overwriting each others’ data. Aggregation code is additional code being included when multiple implementations for one activity are selected and that aggregates the output data of those implementations. Finally, FeatureIDE is used as tool support for the phases domain analysis, domain implementation, requirements analysis, and product derivation.
We implemented the issuing of a special parking permit for craftspersons as a proof-of-concept, which is a common business process in German municipalities. Then, we evaluated our approach and the proof-of-concept by conducting expert interviews which suggest that our proof-of-concept is applicable in many German municipalities and that digitising business processes with varying requirements for activity implementations is a widespread challenge in German municipalities. Nevertheless, in order to test the applicability and practicability of our approach and tool chain, in future work, we want to carry out case studies in which we digitise further business process variants in order to asses the scalability and generalisability to other business processes and domains.
@@
Furthermore, the presented approach has limitations, which will be addressed in future research. It focuses on varying activity implementations and assumes a fixed and immutable control flow of the process model during PAIS Product derivation. Although the expert interviews we conducted during our evaluation revealed that lots of German municipalities already can benefit from our approach, there are business processes whose control flow needs to be customised as well. Future research shall reveal how the control flow may be tailored during PAIS Product derivation. The approaches of base and reference process models might be combined and included into the approach of PAIS Product Lines. Besides, only activities may constitute a variation point, i.e only the implementation of activities may be selected and substituted. Future research shall outline how events may constitute a variation point and how their implementation may be substituted.
Acknowledgement
This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors.
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the authors used Gemini in order to check spelling, grammar, and wording. After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
Appendix A Interview Questions
Introduction of interview participants and introduction to the topic.
  1. 1.
    Do you consider the digitisation of business processes in municipalities important and do you consider it a facilitation of work? Answer options: 1-5 (1: not at all – 5: very much)
  2. 2.
    Can you briefly explain the business process for issuing a special parking permit for craftspersons in your municipality?
  3. 3.
    What information/documents are requested to apply for a special parking permit for craftspersons in your municipality?
    1. Possible answers (others are allowed)
    2. (a)
      Applicant’s name
    3. (b)
      Contact details
    4. (c)
      Vehicle data
      1. i.
        Registration certificate
      2. ii.
        Photos of the vehicle, trunk, and tools
    5. (d)
      Intended use for the special parking permit
    6. (e)
      Business registration (e.g. confirmation by the Chamber of Commerce and Industry or extract from the commercial register)
  4. 4.
    Do you use standard software? Do you use in-house developed software?
    1. Possible answers
    2. (a)
      Standard software
    3. (b)
      In-house developed software
    4. (c)
      PDF/non-digitised
  5. 5.
    Has your business process for issuing a special parking permit for craftspersons changed after digitisation? (Different steps/Different order/Different data) Answer options: 1-5 (1: not at all – 5: very much)
  6. 6.
    Does the digital business process for the special parking permit for craftspersons facilitate work? If so, to what extent? Answer options: 1-5 (1: not at all – 5: very much)
  7. 7.
    Have there been any other improvements, and if so, what are they? Answer options: 1-5 (1: not at all – 5: very much)
  8. 8.
    If you use standard software, to what extent did it have to be customised (input fields/additional or other steps)? Answer options: 1-5 (1: not at all – 5: very much)
  9. 9.
    To what extent does the information system you use meet your expectations for end-to-end process digitisation? Answer options: 1-5 (1: not at all – 5: very much)
  10. 10.
    Are there any features you would like to have that your information system does not offer? And if so, which ones (e.g. a QR code to check the validity of the special parking permit)? Answer options: 1-5 (1: not at all – 5: very much)
At this point in the interview, a demo of our proof-of-concept for the special parking permit for craftspersons is presented, including an explanation of the concepts for customising the business process.
  1. 11.
    Would the presented information system for the special parking permit for craftspersons be applicable in your municipality if it were tailored to your needs using only the presented concepts (customising data, exchanging activity implementations during various binding times)?
  2. 12.
    Does the information system seem easy to learn and intuitive from an applicant’s and municipal clerk’s perspective? Answer options: 1-5 (1: not at all – 5: very much)
  3. 13.
    Do you know of other municipalities that use the special parking permit for craftspersons in this or a different way?
  4. 14.
    Do you know of other municipal business processes that differ between municipalities? If so, which ones and what are the differences?
  5. 15.
    In your opinion, could the concept of the presented information system for the special parking permit for craftspersons be transferred to other municipal business processes (including reasons)?